# Dynamical thermalization, Rayleigh-Jeans condensate, vortexes and wave collapse in quantum chaos fibers and fluid of light

### v2

@@ Abstract
We study analytically and numerically the time evolution of a nonlinear field described by the nonlinear Schrödinger equation in a chaotic $D$-shape billiard. In absence of nonlinearity the system has standard properties of quantum chaos. This model describes a longitudinal light propagation in a multimode D-shape optical fiber and also those in a Kerr nonlinear medium of atomic vapor. We show that, above a certain chaos border of nonlinearity, chaos leads to dynamical thermalization with the Rayleigh-Jeans thermal distribution and the formation of the Rayleigh-Jeans condensate in a vicinity of the ground state accumulating in it about 80-90% of total probability. Certain similarities of this phenomenon with the Fröhlich condensate are discussed. Below the chaos border the dynamics is quasi-integrable corresponding to the Kolmogorov-Arnold-Moser integrability. The evolution to the thermal state is characterized by an unusual entropy time dependence with an increase on short times and later significant decrease when approaching to the steady-state. This behavior is opposite to the Boltzmann H-theorem and is attributed to the formation of Rayleigh-Jeans condensate and presence of two integrals of motion, energy and norm. At a strong focusing nonlinearity we show that the wave collapse can take place even at sufficiently high positive energy being very different from the open space case. Finally for the defocusing case we establish the superfluid regime for vortex dynamics at strong nonlinearity. System parameters for optical fiber experimental studies of these effects are also discussed.
fibers; dynamical chaos; thermalization; wave collapse; vortexes; light superfluidity
@@ I Introduction
In recent years a number of interesting physical phenomena have been established for laser beam propagation in multimode optical fibers (see e.g. overview aplbabin ). Among them there is a phenomenon of self-cleaning of the laser beam when an injected laser power at high modes is transferred to low-energy modes. It was realized that this effect appears due to thermalization among linear fiber modes induced by nonlinear four-wave interactions between modes wabnitz ; fiber1 ; fiber2 ; fiber3 ; fiber4 ; fiber5 . Thus, even thermalization with negative temperatures has recently been observed in such optical fibers fiber5 . It is argued that this thermalization emerges as a result of purely Hamiltonian dynamics without any external thermal bath.
@@
Another important feature is that usually for fibers a nonlinearity is focusing $\beta<0$ and thus, according to the so called Vlasov-Petrishchev-Talanov theorem talanov ; kuznetsov , a wave collapse can take place for strong nonlinearity at finite times. Important implications of such a collapse for Langmuir waves was discussed in zakharov . However, in talanov ; kuznetsov ; zakharov a collapse was discussed in a case of infinite system size. For a D-shape fiber chaos and thermalization in a finite size billiard may lead to an unusual interplay between collapse and dynamical thermalization so that there is a fundamental interest to study the properties of NSE in a D-shape focusing fiber. Here we present results of our studies of such a system considering both cases of negative (focusing) and positive (defocusing) signs of nonlinearity. We show that that the dynamical thermalization with the Rayleigh-Jeans distribution takes place at moderate strength of nonlinearity being above the chaos border. Below this border the dynamics is quasi-integraable corresponding to the KAM theory.
@@
However, NSE describes classical nonlinear fields without second quantization and thus in this case we should have the classical Rayleigh-Jeans (RJ) thermal energy equipartition with chemical potential $\mu$ due to an additional conservation integral of norm or number of particles, see e.g. landau ; zakharovbook :
@@
In the distributions (1), (2) $T$ is a system temperature, $\mu$ is a chemical potential and the total norm of probabilities $\rho_{m}$ on $N$ energy levels $E_{m}$ is preserved and normalized to unity (${\sum^{N}}_{m=1}\rho_{m}=1$; $1\leq m\leq N$). In optical fibers such classical thermal distribution (2) is usually called as Rayleigh-Jeans (RJ) distribution (see e.g. fiber1 ). Indeed, in rmtprl for the NLIRM and related models it was shown that above the chaos border the dynamical chaos at moderate nonlinearity leads to the RJ thermal distribution and not to the BE one. For dynamical systems like NSE we assume that the nonlinear term is relatively weak or moderate and it leads to dynamical chaos for nonlinearity above the chaos border producing thermal distribution (2) over linear eigenenergies (or mode eigenfrequencies) that are not significantly affected by weak nonlinearity.
@@
We argue that the self-cleaning effect aplbabin ; wabnitz is an example of the RJ condensate appearing due to chaos induced by nonlinear interaction of waves in multimode fibers. The emergence of such condensation was also reported for numerical simulations of NSE in two and three dimensional finite systems with quadratic spectrum of waves picozzi1 ; picozzi2 ; picozzi3 ; picozzi2024 . In these works the emergence of RJ distribution was attributed to the Kolmogorov-Zakharov spectra of turbulence zakharovbook . However, as argued above, we attribute the appearance of RJ thermal distribution in these numerical simulations to the degeneracy of linear eigenmode frequencies due to which the KAM theory cannot be applied to these systems and thus chaos appears at very weak nonlinearity strength. For the case of focusing fibers we discuss an interesting interplay between the wave collapse and chaos. We also consider the problem of ultraviolet catastrophe at high energy modes in a chaotic regime.
@@ II Model description
[…]
@@ IV.2 Numerical results
Figure 6: Color map of condensate probability $\rho_{1}$ at corresponding system temperature $T$ and number of system modes $N$ (left panel); right panel shows the same $\rho_{1}$ as in left panel but in the plane $(N,\epsilon_{0})$; $\Delta=1$ obtained in the frame of RJ thermal distribution (2).
@@
[…]
@@
Figure 8: Dependence of the non-condensate probability, $W_{ex}=1-\rho_{1}$, on the total number of modes, $N$, at a fixed total energy of $\epsilon_{0}=2;2.5;3;3.5;4;5;6$. Here, $\rho_{1}$ represents the ground state population. The dashed curve shows the analytical theory given by the average Gumbel distribution with $\rho_{1}=\epsilon_{0}(\log{N}+\gamma)/N$, for $\epsilon_{0}=2$ where $\gamma$ is the Euler-Mascheroni constant (see text).
@@
Figure 9: Color plot of condensate probability $\rho_{1}$. The left panels show plots of $T$ vs. $N$, while the right panels show the distributions of $\epsilon_{0}$ vs. $N$. The top panels display the case of the D-shape billiard, with energies $E_{m}$ of the D-shape fiber described in Section II. The bottom panels show the corresponding plots for the Sinai oscillator, with energies $E_{m}$ from sinaiosc .
@@
Figure 10: Entropy ($S$) as a function of energy ($E$) for Bose-Einstein (BE) distributions (red solid line, Eq. 1) and Rayleigh-Jeans (RJ) distributions (blue dashed line, Eq. 2). The left panels show the simple model with equidistant energy levels, while the right panels are for energy states given by the $D$-billiard. The top panels show the case of $N=100$ states, and the bottom panels for $N=500$ states. For the D-shape billiard the lowest energies are $E_{1}=6.55,E_{2}=15.18,E_{3}=17.8,E_{4}=28.59,E_{5}=28.65$.
@@
In the works rmtprl ; eplbill ; sinaiosc the dependence of entropy on system energy $S(E)$ was used to obtain a distinction between BE (1) and RJ (2) distributions. Thus in Figure 10 we present these two theoretical dependencies for the cases of RJS model of equidistant levels (left panels) and the system with energies of D-shape billiard. The theory shows that there is a significant difference between two dependencies that grows with the increase of total number $N$ of levels in the system that boosts the condensate fraction (see Figure 8). At low energies $E$ the RJ entropy is significantly smaller compared to BE case due to RJ condensation. The results obtained from the NSE dynamics (3) for the dependence $S(E)$ are discussed in the next Section V.
IV.3 Analytical results
@@
To describe analytically the population of eigenstates in the thermodynamic limit of a large number of modes $N$, we assume a microcanonical ensemble for a simplified model with evenly spaced eigen-energies $\epsilon_{n}=n-1$. In this case the total energy is given by:
@@
[…]
@@
[…]
@@
which coincides from the behavior expected from the Rayleigh-Jeans distribution. Thus the condensate fraction $\langle\rho_{1}\rangle$ tends to 100% in the $N\gg 1$ limit. In the context of the time evolution of non-linear Schrödinger equation, this would correspond to a slow depletion eigen-modes at intermediate energies, in favor of a small and continuously decreasing population of modes of higher energy.
It is interesting to know what can limit the condensation in the large $N$ limit in this statistical framework. Including a non-inform eigen-energy energy distribution as well as a weak nonlinearity in the expression of the total energy $\epsilon$ as function of the eigen-level populations $\rho_{n}$ does not seem to stop condensation as long as there is a gap between the ground state and higher excited states. A possible constraint that can limit the condensation fraction is the requirement that the population $\rho_{i}$ do not enter the stability island around the functional ground-state of the nonlinear Schrödinger equation. Since we do not know the equation for the boundary of this stability island we cannot perform realistic simulations, but we checked that simple approximation for the shape of the stability island indeed limit the condensation. For example, if the stability island imposes $\sum_{i=2}^{N}\rho_{i}>\rho^{*}$, then the condensate fraction will converge to $1-\rho^{*}$, with $\rho_{2}\rightarrow\rho^{*}$ in the limit $N\gg 1$.
@@
To summarize, for all systems discussed above we see that the RJ condensate fraction can be high only at low temperature $T\approx\epsilon_{0}/N$ being comparable with a typical energy level spacing when the BE thermal distribution (1) cannot be approximated by the RJ distribution (2). This shows the difference of RJ condensate from the Fröhlich condensate frohlich1 ; frohlich2 existing at high temperature in quantum systems with BE thermal distribution (1).

### v3

@@ Abstract
We study analytically and numerically the time evolution of a nonlinear field described by the nonlinear Schrödinger equation in a chaotic $D$-shape billiard. In absence of nonlinearity the system has standard properties of quantum chaos. This model describes a longitudinal light propagation in a multimode D-shape optical fiber and also those in a Kerr nonlinear medium of atomic vapor. We show that, above a certain chaos border of nonlinearity, chaos leads to dynamical thermalization with the Rayleigh-Jeans thermal distribution and the formation of the Rayleigh-Jeans condensate in a vicinity of the ground state accumulating in it about 80-90% of total probability. Certain similarities of this phenomenon with the Fröhlich condensate are discussed. Below the chaos border the dynamics is quasi-integrable corresponding to the Kolmogorov-Arnold-Moser integrability. We describe also the time evolution during the process of relaxation to the thermal state and the time dependence of quantum von Neumann and classical Boltzmann entropies during this process. At a strong focusing nonlinearity we show that the wave collapse can take place even at sufficiently high positive energy being very different from the open space case. Finally for the defocusing case we establish the superfluid regime for vortex dynamics at strong nonlinearity. System parameters for optical fiber experimental studies of these effects are also discussed.
@@ I Introduction
In recent years a number of interesting physical phenomena have been established for laser beam propagation in multimode optical fibers (see e.g. overview aplbabin ). Among them there is a phenomenon of self-cleaning of the laser beam when an injected laser power at high modes is transferred to low-energy modes. It was realized that this effect appears due to thermalization among linear fiber modes induced by nonlinear four-wave interactions between modes wabnitz ; fiberrev2a ; fiber1 ; fiber2 ; fiber3 ; fiber4 ; christo1 ; fiber5 ; fiberrev2b . Thus, even thermalization with negative temperatures has recently been observed in such optical fibers fiber5 . It is argued that this thermalization emerges as a result of purely Hamiltonian dynamics without any external thermal bath.
@@
Another important feature is that usually for fibers a nonlinearity is focusing $\beta<0$ and thus, according to the so called Vlasov-Petrishchev-Talanov theorem talanov ; kuznetsov , a wave collapse can take place for strong nonlinearity at finite times. Important implications of such a collapse for Langmuir waves was discussed in zakharov . The collapse properties in multimode fibers have been also studied in more recent publications (see e.g. turitsyn ). However, in talanov ; kuznetsov ; zakharov a collapse was discussed in a case of infinite system size. For a D-shape fiber chaos and thermalization in a finite size billiard may lead to an unusual interplay between collapse and dynamical thermalization so that there is a fundamental interest to study the properties of NSE in a D-shape focusing fiber. Here we present results of our studies of such a system considering both cases of negative (focusing) and positive (defocusing) signs of nonlinearity. We show that that the dynamical thermalization with the Rayleigh-Jeans distribution takes place at moderate strength of nonlinearity being above the chaos border. Below this border the dynamics is quasi-integrable corresponding to the KAM theory.
@@
However, NSE describes classical nonlinear fields without second quantization and thus in this case we should have the classical Rayleigh-Jeans (RJ) thermal energy equipartition with chemical potential $\mu$ due to an additional conservation integral of norm or number of particles, see e.g. landau ; zakharovbook ; nazarenko :
@@
In the distributions (1), (2) $T$ is a system temperature, $\mu$ is a chemical potential and the total norm of probabilities $\rho_{m}$ on $N$ energy levels $E_{m}$ is preserved and normalized to unity (${\sum^{N}}_{m=1}\rho_{m}=1$; $1\leq m\leq N$). We use here normalization of all $\rho_{m}$ values to be equal to unity so that $\rho_{m}$ can be considered as a probability at a state $m$, but in a case of many particles $N_{p}$ the value $\rho_{m}$ gives simply a fraction of total number of particle $N_{p}$ in the state $m$. In optical fibers such classical thermal distribution (2) is usually called as Rayleigh-Jeans (RJ) distribution (see e.g. fiber1 ). Indeed, in rmtprl for the NLIRM and related models it was shown that above the chaos border the dynamical chaos at moderate nonlinearity leads to the RJ thermal distribution and not to the BE one. For dynamical systems like NSE we assume that the nonlinear term is relatively weak or moderate and it leads to dynamical chaos for nonlinearity above the chaos border producing thermal distribution (2) over linear eigenenergies (or mode eigenfrequencies) that are not significantly affected by weak nonlinearity.
@@
It is established experimentally and theoretically that the self-cleaning effect aplbabin ; wabnitz ; fiberrev2a ; fiber1 is an example of the RJ condensate appearing at the lowest frequency fiber mode. Here we show that this happens due to Hamiltonian chaos induced by nonlinear interaction of waves in multimode quantum chaos fibers. The emergence of such condensation was also reported for numerical simulations of NSE in two and three dimensional finite systems with quadratic spectrum of waves picozzi1 ; picozzi2 ; picozzi3 ; picozzi2024 . In these works the emergence of RJ distribution was attributed to the Kolmogorov-Zakharov spectra of turbulence zakharovbook ; nazarenko . However, as argued above, we attribute the appearance of RJ thermal distribution in these numerical simulations to the degeneracy of linear eigenmode frequencies due to which the KAM theory cannot be applied to these systems and thus chaos appears at very weak nonlinearity strength. For the case of focusing fibers we discuss an interesting interplay between the wave collapse and chaos. We also consider the problem of ultraviolet catastrophe at high energy modes in a chaotic regime.
@@ II Model description
where we consider $\hbar=1$, $m=0.5$, $\vec{r}$ is a vector in $(x,y)$ plane. The form of D-shape fiber billiard is shown in Fig. 1. The time variable $t$ corresponds to $z$ direction of a beam propagation along fiber (see e.g. Eq.(1) in aplbabin ), $\psi$ represents a field inside fiber and $\beta$ is a field strength coefficient of nonlinearity. This equation has two integrals of motion being the norm $\int|\psi(\vec{r},t)|^{2}dxdy=1$ and energy       $E=<\psi(t)|\hat{H_{0}}|\psi(t)>+\beta|\psi(t)|^{4}/2$ where $\hat{H}_{0}$ is the operator of linear Hamiltonian at $\beta=0$.
@@ IV.2 Numerical results
Figure 6: Color map of condensate probability $\rho_{1}$ at corresponding system temperature $T$ and number of system modes NN (left panel); right panel shows the same $\rho_{1}$ as in left panel but in the plane $(N,\epsilon_{0})$; $\Delta=1$ obtained in the frame of RJ thermal distribution (2).
@@
Figure 7: Thermal RJ distribution $\rho_{m}$ (2) for various mode energies $E_{m}$ at $N=20,50,100,300$ and initial total energy $\epsilon_{0}\ 2.5$; $\Delta=1$; data is presented in log-log scale.
@@
Figure 8: Dependence of the non-condensate probability, $W_{ex}=1-\rho_{1}$, on the total number of modes, NN, at a fixed total energy of $\epsilon_{0}=2;2.5;3;3.5;4;5;6$. Here, $\rho_{1}$ represents the ground state population. The dashed curve shows the analytical theory given by the average Gumbel distribution with $\rho_{1}=\epsilon_{0}(\log{N}+\gamma)/N$, for $\epsilon_{0}=2$ where $\gamma$ is the Euler-Mascheroni constant (see text).
@@
Figure 9: Color plot of condensate probability $\rho_{1}$. The left panels show plots of $T$ vs. NN, while the right panels show the distributions of $\epsilon_{0}$ vs. NN. The top panels display the case of the D-shape billiard, with energies $E_{m}$ of the D-shape fiber described in Section II. The bottom panels show the corresponding plots for the Sinai oscillator, with energies $E_{m}$ from sinaiosc .
@@
IV.3 Classical Boltzmann entropy $S_{B}$ and quantum von Neumann entropy $S_{q}$
The classical Boltzmann entropy $S_{B}=\ln\Gamma$ is determined by a phase volume $\Gamma$ of microscopic states at a given energy boltzmann1 ; mayer ; landau . For multimode fiber with a finite number of modes, as e.g. for the above RJS model, it was shown picozzi2 ; picozzi3 ; picozzi2024 ; christo2021 that
$S_{B}=\sum_{m}\ln(\rho_{m})+C\;\;,$ (7)
where $C$ is some constant that can make entropy to be positive, but we take $C$ to be zero. We note that the total number of particles, or norm, and the total energy $E=\epsilon_{0}$ are preserved in our system and thus these two integrals determine the functions $T(E)$ and $\mu(E)$ entering in RJ distribution (2). From the above expression for $S_{B}$ and RJ thermal distribution (2) it follows that $S_{B}$ satisfies the standard thermodynamic relation landau ${dS_{B}}/dE=1/T$ where $E$ is the total system energy on which depends temperature $T(E)$ and chemical potential $\mu(E)$ from the expression for $\rho_{m}$ (2). Indeed, the expression for ${dS_{B}}/dE$ directly follows from (2), (4) that gives: $dS_{B}/dE=\sum_{m}(d\rho_{m}/dE)/\rho_{m}=(NdT/dE)/T+(d\mu/dE)/T=1/T$. We note that the relation $dS_{B}/dE=1/T$ works for any sequence of level energies $E_{m}$.
The expression (7) for $S_{B}$ can be used during the time evolution of the initial distribution $\rho_{m}(t)$ over modes $m$ at $t=0$ and higher $t$, so that we have time dependence $S_{B}(t)=\sum_{m}\ln\ (\rho_{m}(t))$. In picozzi2 ; picozzi3 it was shown that for a fiber with a finite number of modes $N$ nonlinearity leads to a monotonic growth of $S_{B}(t)$ with time in agreement with the Boltzmann H-theorem boltzmann1 ; mayer ; landau .
@@ IV.3 Classical Boltzmann entropy $S_{B}$ and quantum von Neumann entropy $S_{q}$
Figure 10: Quantum von Neumann entropy $S_{q}(E)$ as a function of energy ($E$) for Bose-Einstein (BE) distributions (red solid line, Eq. 1) and Rayleigh-Jeans (RJ) distributions (blue dashed line, Eq. 2). The left panels show the RJS model with equidistant energy levels, while the right panels are for energy states given by the $D$-billiard. The top panels show the case of $N=100$ states, and the bottom panels for $N=500$ states. For the D-shape billiard the lowest energies are $E_{1}=6.55,E_{2}=15.18,E_{3}=17.8,E_{4}=28.59,E_{5}=28.65$.
At $\beta=0$ we have a quantum system (3) and it is tenting to assume that the small or moderate nonlinear $\beta$-term acts as a some kind of dynamical nonlinear thermostat eplbill ; sinaiosc . In such a case the distribution $\rho_{m}$ of probability over billiard eigenmodes would be described by the BE thermal distribution (1) with the system quantum Neumann entropy neumann (see also mayer ; landau ; petz ):
$S_{q}=-\sum_{m}\rho_{m}\ln(\rho_{m})\;\;.$ (8)
However, in the system (3) there is no second quantization and in fact we have a nonlinear dynamics of classical field where the thermal distribution of $\rho_{m}$ is given by the RJ relation (2) as it was shown for nonlinear perturbation of RMT rmtprl . Thus in our classical system (3) with the RJ distribution (2) we have the Boltzmann entropy $S_{B}$ that satisfies the thermal relation $dS_{B}/DE=1/T$ while for the quantum entropy $S_{q}$ (8) and RJ distribution (2) we have $dS_{q}/dE\neq 1/T$.
However, it is still useful to study the properties of quantum entropy $S_{q}$ since it gives the effective number of populated eigenmodes $\Delta m_{eff}=\exp(S_{q})$. The time dependence $S_{q}(t)$ is given by the relation (8) where one should take $\rho_{m}(t)$ at a given moment of time so that $S_{q}(t)=-\sum_{m}\rho_{m}(t)\ln(\rho_{m}(t))$ .
@@
In the works rmtprl ; eplbill ; sinaiosc the dependence of quantum entropy on system energy $S_{q}(E)$ was used to obtain a distinction between BE (1) and RJ (2) distributions. Thus in Figure 10 we present these two theoretical dependencies for the cases of RJS model of equidistant levels (left panels) and the system with energies of D-shape billiard. The theory shows that there is a significant difference between two dependencies that grows with the increase of total number $N$ of levels in the system that boosts the condensate fraction (see Figure 8). At low energies $E$ the RJ entropy is significantly smaller compared to the BE case due to RJ condensation.
The results obtained from the NSE dynamics (3) determine the dependences $S_{q}(E,t)$ and $S_{B}(E,t)$ which are discussed in the next Section V.
IV.4 Analytical results
@@
To describe analytically the population of eigenstates in the thermodynamic limit of a large number of modes $N$, we assume a microcanonical ensemble for a simplified model with evenly spaced eigenenergies $\epsilon_{n}=n-1$. In this case the total energy is given by:
@@
${\hat{s}}_{N}(\beta,\mu)$ $=$ $\int_{0}^{\infty}d\epsilon e^{-\beta\epsilon}\int_{0}^{\infty}dpe^{-\mu p}s_{N}(\epsilon,p)$ (12)
@@
$s_{N}(\epsilon,p)=\frac{N}{2\pi i}\int d\beta\exp\left(\epsilon\beta+(N-1)\log\frac{1-e^{-p\beta}}{\beta}\right)$ (17)
@@
which coincides from the behavior expected from the Rayleigh-Jeans distribution. Thus the condensate fraction $\langle\rho_{1}\rangle$ tends to 100% in the $N\gg 1$ limit. In the context of the time evolution of non-linear Schrödinger equation, this would correspond to a slow depletion eigenmodes at intermediate energies, in favor of a small and continuously decreasing population of modes of higher energy.
It is interesting to know what can limit the condensation in the large $N$ limit in this statistical framework. Including a non-inform eigenenergy energy distribution as well as a weak nonlinearity in the expression of the total energy $\epsilon$ as function of the eigenlevel populations $\rho_{n}$ does not seem to stop condensation as long as there is a gap between the ground state and higher excited states. A possible constraint that can limit the condensation fraction is the requirement that the population $\rho_{i}$ do not enter the stability island around the functional ground-state of the nonlinear Schrödinger equation. Since we do not know the equation for the boundary of this stability island we cannot perform realistic simulations, but we checked that simple approximation for the shape of the stability island indeed limit the condensation. For example, if the stability island imposes $\sum_{i=2}^{N}\rho_{i}>\rho^{*}$, then the condensate fraction will converge to $1-\rho^{*}$, with $\rho_{2}\rightarrow\rho^{*}$ in the limit $N\gg 1$.
@@
IV.5 RJ condensate and Fröhlich condensate
To summarize, for all systems discussed above we see that the RJ condensate fraction can be high only at low temperature $T\approx\epsilon_{0}/N$ being comparable with a typical energy level spacing (see Figures 69). In contrast to this condition $T\ll\Delta\sim E_{m}$ the BE thermal distribution (1) can be approximated by the RJ distribution (2) only at high temperature $T\gg E_{m}$. Thus it is not possible to start to approximate the BE distribution (1) of a quantum system at high temperature, obtain the RJ distribution (2) and then to have RJ condensate since it appears only at small temperature $T\sim T_{c}\sim\Delta$ when the approximation of BE by RJ distribution is not valid. In fact such a logical scheme was used by Fröhlich in frohlich1 ; frohlich2 . Thus Fröhlich in fact follows this RJ approximation for BE at high temperatures and then finds that RJ distribution has the condensate phase (that formally is not logically correct since RJ condensate exists only at low temperature while transition from BE to RJ distribution is valid only at high temperature). However, at the same time Fröhlich assumes that there is external energy pumping of the system which is absorbed by certain dissipative processes in the system so that the conditions of possible realization of Fröhlich condensate are significantly complicated by the system couplings with external sources while the RJ condensate appears in a completely isolated system. This system is a Hamiltonian conservative system described by classical nonlinear fields so that from the begging the thermal distribution of this classical system is given by the RJ distribution (2) without any links with the quantum BE distribution (1).